# DocFinQA: A Long-Context Financial Reasoning Dataset

### v2

@@ Abstract
For large language models (LLMs) to be effective in the financial domain – where each decision can have a significant impact – it is necessary to investigate realistic tasks and data. Financial professionals often interact with documents that are hundreds of pages long, but most financial research datasets only deal with short excerpts from these documents. To address this, we introduce a long-document financial QA task. We augment 7,437 questions from the existing FinQA dataset with the full-document context, extending the average context length from under 700 words in FinQA to 123k words in DocFinQA. We conduct extensive experiments over retrieval-based QA pipelines and long-context language models. DocFinQA proves a significant challenge for even state-of-the-art systems. We also provide a case-study on the longest documents in DocFinQA and find that models particularly struggle on these documents. Addressing these challenges may have a wide reaching impact across applications where specificity and long-range contexts are critical, like gene sequences and legal document contract analysis. The data and code is publicly accessible at github.com/anonymous.
@@ 1 Introduction
The frequent need to reason over large volumes of textual and tabular data makes financial analysis particularly challenging for LLMs Azzi et al. (2019). Existing work on automating financial numerical reasoning focuses on unrealistically specific document snippets Chen et al. (2021); Zhu et al. (2021). Datasets are often limited to pre-selected document sections, failing to reflect the broader and more realistic scenarios faced by analysts Masson and Montariol (2020). Financial professionals often sift through hundreds of pages per document, requiring deep understanding of both the content and structure to effectively navigate and extract pertinent information. Current long-document QA datasets such as NarrativeQA Kočiský et al. (2018) do not test the quantitative reasoning skills needed in the financial domain.
In this work, we introduce DocFinQA, a long-document financial question answering task. We extend the FinQA dataset of expert annotated questions and answers Chen et al. (2021) with full Securities and Exchange Commission (SEC) reports. This results in a significantly longer context in the DocFinQA dataset – by a factor of 175 – than the FinQA dataset. The resulting long-document QA task offers a more realistic evaluation of a model’s reasoning capabilities over financial documents. In line with recent work on program synthesis for financial QA Koncel-Kedziorski et al. (2023), the questions in DocFinQA are annotated with Python programs to generate the answers, allowing for training and evaluating program synthesis models for use in realistic financial workflows.
@@ 2 Related Work
Prior studies in financial question answering focus on non-numerical reasoning Day and Lee (2016); Jørgensen et al. (2023); Maia et al. (2018). Short-context grounded numerical reasoning tasks were introduced with datasets such as FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021). Recently, understanding long documents has attracted more attention for tasks involving events Yang et al. (2018), table of contents Bentabet et al. (2020), and causal relations Mariko et al. (2022). However, to the best of our knowledge, this is the first attempt to address financial numerical QA by grounded in long documents with upwards of hundreds of pages of context for each question.
Long-document QA has been studied in NLP with the introduction of datasets such as SearchQA Dunn et al. (2017), NarrativeQA Kočiský et al. (2018), QuALITY Pang et al. (2022), and PDFTriage Saad-Falcon et al. (2023). Due to the limited context size of LLMs, retrieval-based models are commonly used to filter irrelevant text Izacard et al. (2022); Lewis et al. (2020). Recently, advances in attention mechanisms Beltagy et al. (2020); Dao et al. (2022) and positional embeddings Press et al. (2021); Su et al. (2023) allow for end-to-end grounded QA with context windows of more than 100k tokens. However, these methods suffer from loss of important context Zhang et al. (2023) and often fail to make full use of longer inputs Liu et al. (2023). Our work studies the intersection of numerical reasoning and long-document processing, and our results demonstrate that there is still ample room for improvement in this domain.
Dataset #Docs #QAs #Words Pag Num Tab
@@
Table 1: Comparison of DocFinQA and existing Finance QA and Long Document QA dataset. DocFinQA includes multi-page documents rich with both numeric data and tables.
3 DocFinQA Dataset
Our dataset is an extension of the FinQA dataset. (See Table 5 for an example of FinQA). FinQA was created by finance experts from S&P 500 companies’ annual financial report (aka., 10-K). They used full documents when creating the questions (e.g., 100s of pages), but FinQA only includes the curated context necessary to produce correct answers (e.g., a paragraph and/or table of statistics).
To recreate the original problem, we retrieve the full SEC filing for each question. Following Koncel-Kedziorski et al. (2023), we require the model to generate Python code to produce an answer instead of directly generating the numeric answer. The generated Python code illustrates what information from the context is used, along with what arithmetic operations are performed.
Dataset Representation: Each question in FinQA is a triplet $(c^{golden},q,a)$ composed of a golden context $c^{golden}$, a question $q$, and an answer $a$ written in human language. We extend the dataset in two ways: (1) context $c^{golden}$ is extended to the full document context $D$; and (2) we added a Python program $p$ that produces the answer $a$. Each final sample in DocFinQA is a quartet $(D,q,p,a)$. See Appendix M for an example.
Filings Collection: For each question of the FinQA dataset, we identify the corresponding SEC filing from which it was originally created. We retrieved the filing in HTML/XML format from SEC’s EDGAR service and parsed text and table into clean markdown format text Wang et al. (2023). The collection and parsing processes are presented in more detail in Appendix F and Appendix G, respectively. Figure 2 shows the distribution of document lengths in DocFinQA.
Figure 2: Histogram of document length (#words) in DocFinQA dataset with dash line representing the average length of the documents. Purple line depicts the proportion of documents where the question context is within the current number of words.
@@ 3 DocFinQA Dataset
Chunking and Alignment: To study retrieval-based QA systems, we split each document $D$ into a set of chunks $C=\{c_{1},\cdots,c_{n}\}$. Each chunk contains 2,750 characters ($\sim\!509$ tokens) with a 20% overlap to avoid missing context at the edges. To compute the performance, we identify the best context chunk, $c^{*}$, from the chunk set $C$ associated with each document $D$ that includes the information to answer question $q$. Since FinQA already provides $c^{golden}$, we compute a pair-wise score $(c_{i},c^{golden})$, for all chunks, $c_{i}\in C$, including the golden chunk. We found that four-gram based similarity score offers the sharpest matching signal among tri-gram, four-gram, and fuzzy matching. The chunk with the highest score is selected as the target context chunk for retrieval. We verify that this process results in good $c^{*}$ chunks through manual inspection and by substituting $c^{*}$ for $c^{golden}$ in a few-shot QA evaluation with GPT-3.5. The model scores slightly higher when using $c^{*}$ rather than $c^{golden}$ (67.5% vs 67.3% accuracy respectively), suggesting that little context is lost.
@@
Code Generation: The FinQA dataset provides solutions in a “program” syntax that, when executed, yields the answer (e.g., in Figure 1 the solution is divide(102400, 619314). However, this derivation does not provide meaningful context of what is being calculated. In our running example, 102400 is not semantically grounded to the document. Koncel-Kedziorski et al. (2023) augments FinQA with readable Python code (including named variables like, dec_shares = 102_400) that can be executed to derive the answer, providing a layer of interpretability. Thus, we use the code-enhanced version of FinQA to construct DocFinQA (See Appendix I).
@@
Statistics: The resultant DocFinQA dataset comprises 5,735 training, 780 development, and 922 test samples, derived from 801 unique SEC filings. Table 1 shows the statistics and characteristics of DocFinQA in comparison with other finance numerical reasoning and long-document QA datasets.
Due to the limited availability of complete SEC filings (refer Appendix F) and imperfections in the code generation process, DocFinQA encompasses a subset of FinQA questions. We analyze the distribution of the question types of FinQA and DocFinQA datasets (See Appendix C). DocFinQA remains a representation of all major question types from the original FinQA dataset.
4 Retrieval-based QA Evaluation
Retrieval Task We test three models for context retrieval: ColBERT (ColB)Khattab and Zaharia (2020), Sentence-BERT (SentB) Reimers and Gurevych (2019), and OpenAI’s Ada Greene et al. (2022). Further, we finetune the ColBERT model (FT ColB) on the training set of DocFinQA to evaluate an in-domain model. We test a matching-based model, BM25 Robertson et al. (1995), but observe poor performance (see Appendix K).
To retrieve context for a question $q$ over chunk set $C$, we encode both $q$ and $C$ with the encoding models mentioned above. This results in an embedding for the question $v_{q}$ and chunk embeddings $V_{C}=\{v_{c_{i}}|c_{i}\in C\}$. We compute the cosine similarity between $v_{q}$ and each vector in $V_{C}$ to retrieve the top-k most similar chunks. We evaluate these models using hit rate (HR@k) on the test set of DocFinQA using the target $c^{*}$. Results are shown in Figure 3. The FT ColB yields the highest HR, followed by ColB. FT ColB yields an average improvement of 91% HR over SentB and obtains a 0.35 (HR@1) and 0.55 (HR@3).
Question Answering Task We formulate the QA task as a few-shot in-context learning task Brown et al. (2020). For each in-context example, we only provide the gold chunk and the correct answer. For the actual query we provide $k$ chunks. A sample prompt template is given in Figure 10.
Table 2 reports the performance of 11 state-of-the art models. Larger models outperform smaller models (e.g., MPT 30B vs MPT 7B). Models trained on code yield higher accuracy than non-code models (e.g., CodeLlama vs Llama). Models with additional supervised finetuning (e.g., LLama 2/7B+SFT) and instruction tuning (e.g., GPT-3.5) are among the best examined. Notably, Mistral 7B outperforms several larger models, although it lags behind Llama 2/7B+SFT and GPT-3.5.
Model/Size ColB SentB ADA FT ColB
Falcon/7B   2.3   0.3   1.2   1.8
MPT/7B   4.6   2.7   3.8   4.8
MPT/30B 17.3 11.1 12.1 18.1
Llama 2/7B 13.5 8.9 11.1 13.5
Llama 2/13B 18.7 12.7 14.9 19.1
CodeLlama/7B 15.6 12.2 15.2 16.8
CodeLlama/13B 19.1 13.8 18.8 21.0
Mistral/7B 23.2 14.9 21.5 25.0
Llama 2/7B+SFT 32.9 24.8 34.3 36.1
GPT-3.5/- 41.6 33.8 36.4 42.6
Table 2: Performance on DocFinQA test set. For each row, the best performance among all retrieval models is in bold. The fewshot setting is selected based on the best performance on the dev set (See Appendix L).
5 Case Study w/ 100K+ Token Documents
Recent LLMs can handle context lengths of 128K tokens, but more than 40% of the documents in DocFinQA remain unanswerable even at this content length (see Figure 2). Here, we evaluate performance on a test subsample of 200111The test subsample is limited to 200 documents due to the monetary and temporal costs of human evaluation and GPT4. long documents, each of which has 100K or more tokens.
We explore two retrieval-free options. System 2 Attention (S2A) extracts relevant information from each 100K-token chunk of a document before answering the question using the combined extracted information as context  (Weston and Sukhbaatar, 2023). The Iterative method produces the output program iteratively as the LLM processes each 100k section of the document. A temporary answer program (initially “None”) is input with each section to the LLM. We also report the performance of the best retrieval-based model (Retrieval) based on the experiment in Section 4. The human evaluation we conducted on these 200 questions highlights the challenging nature of this dataset. [^2] Experienced but non-expert human participants. See Appendix A for details.
Model/Size + Method w/ Retrieval Test Subsample
Human No 41.0
Mistral/7B + Iterative No 11.5
Mistral/7B + S2A No 15.5
Mistral/7B + Retrieval Yes 20.0
GPT-4 + Iterative No 20.0
GPT-4 + S2A No 23.0
GPT-4 + Retrieval Yes 47.5
Non-expert human performance on DocFinQA is lower than FinQA (41% versus 50.7%). This can be attributed to the difficulty of also finding the golden page, compared to the golden page being given in FinQA. Notably, the expert performance reported in FinQA is 91.2%.
Still, non-expert human performance is double that of retrieval-free GPT-4 on these long documents, and roughly triple that of retrieval-free Mistral. Secondly, Iterative performed worse than S2A for both GPT-4 and Mistral with a reduced accuracy of 3% and 4%, respectively. Third, with retrieval, both Mistral and GPT-4 outperform their retrieval-free counterparts, with the assisted GPT-4 now on par with the human cohort. Together, these results highlight that DocFinQA is an effective and difficult test for long-document QA, and that there is still room for significant improvement in this domain. For instance, further exploration into methods that combine information across multiple calls to a document-processing LLM is warranted.
6 Conclusion
This paper introduces a realistic document-level question answering dataset over financial reports. Each question includes a full financial report (averaging 123K words), a far greater challenge than previous work that hones in on pre-specified content. Our findings reveal that this more realistic setting presents a significantly more difficult challenge, thereby opening new avenues for research in quantitative financial question answering.
7 Limitation
This work introduced an extension of the existing FinQA dataset. Due to limited human resources, we only validate the test set while the training and the development set were not fully validated. As a result, we can not make any claim of bias and question quality in the not-yet-validated data points offered in this paper. Additionally, as discussed in section 2, the code provided in this work was generated by WizardCoder LLMs. Our assumption is that the code is correct if it produces correct or approximately close to the golden answer. This method may generate both false positive codes (the code that generate correct answer with incorrect rationales) and false negative codes (the correct code that fail the approximation test).
8 Broader Impact and Ethical Considerations
We do not foresee any considerable risks associated with our work given that it is an extension of a publicly available documents and dataset. To uphold transparency, the paper provides detailed documentation of the dataset creation process, including the sources of data and annotation details. Our dataset serves as a resource to underscore the need for more long context oriented benchmarks both within and outside the financial domain and does not intend to criticize any one or more LLMs.
The annotation in this work is done automatically, so no crowd-source or contract annotators were hired throughout the process. The human evaluation in this study were done by full-time paid coworkers known to the authors.
Appendix A Human Evaluation Setting
Appendix B Retrieval Performance
Figure 3: Hit rate of retrieval models.
Appendix C Impact of Data Selection
@@
DocFinQA was subsampled from the full FinQA dataset based on the correctness of the code produced by an open-source LLM. This process may potentially filter out a collection of question types that were not answered by the LLM due to its limited capability. This section investigates the impact of this process by comparing the distribution of the question types in FinQA and DocFinQA. To do this, we show the distribution of questions grouped by their first 2 non-stop words in the Figure 4. The most important observation is that, overall, the distribution of the question set in DocFinQA and FinQA are very similar. There are no major groups being filtered out by our data selection process. The dominant questions (above 1% in FinQA) remains dominant and no major impact on the percentages of those questions is observed. The mid group (above 0.2% in FinQA) question sets see a mixed effect. A large portion of these questions see an increase in percentage while some experience significant loss (e.g., “what percentual” and “what decrease”). Lastly, the long tail group ( under 0.2% in FinQA) either remains the same (e.g., “percent total” and “what greatest”) or is completely wiped out due to small population (e.g., “was average”, and “what return”).
Figure 4: Distribution of question grouped by question types in the original FinQA and DocFinQA. The x-axis (percentage) is presented in log-scale to magnify the differences between the two sets.
Appendix D Golden Chunk Position
Figure 5 shows the distribution of the position golden chunk with the documents. We see that most of the golden chunks appear within the first 250 chunks (approximately 125K tokens which can be fed into the newest generative models). Nonetheless, there are a substantial number of questions that the golden chunk appear beyond this threshold.
Figure 5: Histogram of position of the FinQA context in the original SEC filing that is split into chunks of size 2750.
Appendix E Model Details
Appendix F SEC Filing Collection
Downloading the right 10-K filing from SEC begins with identifying the company code from the company ticker symbol. For example, C/2017/page_328.pdf-1 in FinQA maps to the CITIGROUP INC with company code 831001. This mapping is obtained from the official file released by SEC which can be found here https://www.sec.gov/file/company-tickers. We automatically generate a URL using the company code obtained. From the SEC website, either filings are downloaded as TXT, HTML or XBRL using the generated URL. At this stage, approximately 6.5% (or 543) data points which is corresponds to approximately 9.4% (or 17) documents were dropped, either due to lack of mapping or nonavailability of older documents. Further, conversion of the downloaded files to PDF caused a loss of 117 data points (19 unique documents) due to formatting issues.
Appendix G Parsing SEC Filings
We explore different methods for parsing SEC filings consisting of HTML and XML markup into text and markdown tables for use in our QA systems. To evaluate parsing strategies, we measure HR@k when searching for the gold chunk among all document chunks for a single document using the FinQA question as the search query. Queries and document chunks are encoded with OpenAI’s ADA model. We compare BeautifulSoup, a standard library for manipulating HTML and XML formatted data, and Kensho Extract, a finance specific text and table extraction model. [^3] Passing the raw HTML/XML to the language model produces near-zero performance.
Figure 6 shows the performance of these two methods. We see that the finance specific models used in Kensho Extract result in better downstream performance of this retrieval setup compared to Beautiful Soup. Qualitative analysis of the different parsers reveals that Kensho Extract is better at structuring the tables used in financial documents, resulting in better readability which seems to extend to the encodings.
Appendix H ColBERT Finetuning
We finetune the original ColBERT v1 model on the train set of DocFinQA. For each data point we perform chunking and alignment to generate one golden chunk and $n-1$ negative chunks. For training, we generate a list of tuples (qid, pid+, pid-), where qid refers to the question, pid+ refers to the golden chunk and pid- refers to each of the negative chunks in that document. We train the model for a total of 3 epochs and store the checkpoints at the end of each epoch. The hit rate of the Finetuned ColBERT model after each epoch on the development set is shown in Figure 7. We observe that after the first epoch, additional finetuning does not show any performance improvement. The Finetuned ColBERT model referred to in this study thus uses the weights after the first epoch of training.
Appendix I Code Conversion
Figure 8 shows the steps of converting (a) derivation of the result in FinQA into (b) dummy Python code with dummy variable names , and finally transform it to (c) a meaningful Python program in DocFinQA following the work by Koncel-Kedziorski et al. (2023).
(a) $subtract(34.8,1.2),divide(\#0,34.8)$ (b) $a=34.8-1.2$ $b=a/34.8$ $c=b*100$ (c) $payments\_decrease=34.8-1.2$ $change=payments\_decrease/34.8$ $answer=change*100$
Appendix J Few-shot Settings
Due to the limited context length of the LLMs, the number of few-shot demonstrations and the number of chunks fed into the In-Context Learning must be optimized. We explore 3 settings of number of few-shot examples and 4 settings of number of chunks used as context in the query. Figure 9 shows the performance of these settings in with retrieval and answered by LLama 13B and CodeLLaMa 13B on the development set. We see that a higher number of few-shot example (numshot=3) yield consistent better performance compared to a lower one (numshot=1).
Figure 10 shows a the prompt template with in-context learning that we used.
Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {golden chunk} Question: {question} Python Program: {program} Answer: {answer} Context: {first chunk} {second chunk} {third chunk} Question: {question} Python Program:
Appendix K Detailed performance of retrieval method
Figure 11 shows a pilot study comparing dense-retrieval with OpenAI ADA and Sentence BERT versus sparse retrieval (BM 25) on the development set. We can clearly see that dense retrieval model offer a much higher hit ratio.
Appendix L Detailed performance on dev set
Table 4 reports the full performance on the development set with four retrieval models and three few-shot settings. This results in a total of twelve unique configurations. For our analysis we employ ten state-of-the-art LLMs and a finetuned version of LLama2 7B. For both finetuned and pretrained models we use greedy decoding whenever applicable. For hyperparameter tuning, we carry out the extraction task on the development set of DocFinQA. One trend noted was that all generic LLMs showed higher accuracy with shorter context and more fewshot example i.e. top chunk with 3 shot. While the code based LLMs such as Starcoder and CodeLLama showed higher accuracy with longer context i.e top 3 chunks with 3 shot. This trend is also depicted in Figure 9.

### v3

@@ Abstract
For large language models (LLMs) to be effective in the financial domain – where each decision can have a significant impact – it is necessary to investigate realistic tasks and data. Financial professionals often interact with documents spanning hundreds of pages, but most financial research datasets only deal with short excerpts from these documents. To address this, we introduce a long-document financial QA task. We augment 7,437 questions from the existing FinQA dataset with full-document context, extending the average context length from under 700 words in FinQA to 123k words in DocFinQA. We conduct extensive experiments over retrieval-based QA pipelines and long-context language models. Based on our experiments, DocFinQA proves a significant challenge for even state-of-the-art systems. We also provide a case study on a subset of the longest documents in DocFinQA and find that models particularly struggle with these documents. Addressing these challenges may have a wide-reaching impact across applications where specificity and long-range contexts are critical, like gene sequences and legal document contract analysis. DocFinQA dataset is publicly accessible [^1] https://huggingface.co/datasets/kensho/DocFinQA.
@@ 1 Introduction
The frequent need to reason over large volumes of textual and tabular data makes financial analysis particularly challenging for LLMs Azzi et al. (2019). Existing work on automating financial numerical reasoning focuses on unrealistically specific document snippets Chen et al. (2021); Zhu et al. (2021). Datasets are often limited to pre-selected document sections, failing to reflect the broader and more realistic scenarios faced by analysts Masson and Montariol (2020). Financial professionals usually sift through hundreds of pages per document, requiring a deep understanding of both content and structure to navigate and extract pertinent information effectively. Current long-document QA datasets such as NarrativeQA Kočiský et al. (2018) do not test the quantitative reasoning skills needed in the financial domain.
@@
In this work, we introduce DocFinQA, a long-document financial question-answering task. We extend the FinQA dataset of expert annotated questions and answers Chen et al. (2021) with full Securities and Exchange Commission (SEC) reports. This results in a significantly longer context in the DocFinQA dataset – by a factor of 175 – than the FinQA dataset. Additionally, we manually verified and annotated questions of the test set. The resulting long-document QA task offers a more realistic evaluation of a model’s reasoning capabilities over financial documents. In line with recent work on program synthesis for financial QA Koncel-Kedziorski et al. (2023), the questions in DocFinQA are appended with Python programs to generate the answers, allowing for training and evaluating program synthesis models for use in realistic financial workflows.
@@ 2 Related Work
Prior studies in financial question answering focus on non-numerical reasoning Day and Lee (2016); Jørgensen et al. (2023); Maia et al. (2018). Short-context grounded numerical reasoning tasks were introduced with datasets such as FinQA Chen et al. (2021) and TAT-QA Zhu et al. (2021). Recently, understanding long documents has attracted more attention for tasks involving events Yang et al. (2018), table of contents Bentabet et al. (2020), and causal relations Mariko et al. (2022). However, to the best of our knowledge, this is the first attempt to address financial numerical QA grounded in long documents with upwards of hundreds of pages of context for each question.
Comparison of DocFinQA and existing Finance QA and Long Document QA Long-document QA has been studied in NLP with the introduction of datasets such as SearchQA Dunn et al. (2017), NarrativeQA Kočiský et al. (2018), QuALITY Pang et al. (2022), and PDFTriage Saad-Falcon et al. (2023). Due to the limited context size of LLMs, retrieval-based models are commonly used to filter irrelevant text Izacard et al. (2022); Lewis et al. (2020). Recently, advances in attention mechanisms Beltagy et al. (2020); Dao et al. (2022) and positional embeddings Press et al. (2021); Su et al. (2023) allow for end-to-end grounded QA with context windows of more than 100k tokens. However, these methods suffer from loss of important context Zhang et al. (2023) and often fail to make full use of longer inputs Liu et al. (2023). Our work studies the intersection of numerical reasoning and long-document processing, and our results demonstrate that there is still ample room for improvement in this domain.
3 DocFinQA Dataset
Dataset Representation Each question in FinQA is a triplet $(c^{golden},q,a)$ composed of a golden context $c^{golden}$, a question $q$, and an answer $a$ written in human language. An example of FinQA is shown in Table 5 (See Appendix A). We extend the dataset in two ways: (1) context $c^{golden}$ is extended to the full document context $D$, and (2) we added a Python program $p$ that produces the answer $a$. Each final sample in DocFinQA is a quartet $(D,q,p,a)$. An example of DocFinQA is shown in Table A (See Appendix A).
Filings Collection: For each question of the FinQA dataset, we identify the corresponding SEC filing from which it was created. We retrieve the filing in HTML/XML format from SEC’s EDGAR service and parse the text and table into clean markdown format Wang et al. (2023). The collection and parsing processes are presented in more detail in Appendix A and Appendix B, respectively. Figure 2 shows the distribution of document lengths in DocFinQA.
Figure 2: Histogram of document length (#words) in DocFinQA dataset with dash line representing the average length of the documents. The purple line depicts the proportion of documents where the question context is within the current number of words.
Dataset #Docs #QAs #Words Multi-page Numeric Tabular
@@ 3 DocFinQA Dataset
Table 1: Comparison of DocFinQA and existing Finance QA and Long Document QA dataset. DocFinQA includes multi-page documents with both numeric and tabular data.
@@
Chunking and Alignment: To study retrieval-based QA systems, we split each document $D$ into a set of chunks $C=\{c_{1},\cdots,c_{n}\}$. Each chunk consists of 2,750 characters ($\sim\!509$ tokens) with a 20% overlap to avoid missing context at the edges. To compute the performance, we identify the best context chunk, $c^{\bigstar}$, from the chunk set $C$ associated with each document $D$ that includes the information to answer question $q$. Since FinQA already provides $c^{golden}$, we compute a pair-wise score $(c_{i},c^{golden})$, for all chunks, $c_{i}\in C$, including the golden chunk. We find that four-gram-based similarity score offers the sharpest matching signal among tri-gram, four-gram, and fuzzy matching. The chunk with the highest score is selected as the target context chunk for retrieval. We verify that this process results in good $c^{\bigstar}$ chunks through manual inspection and by substituting $c^{\bigstar}$ for $c^{golden}$ in a few-shot QA evaluation with GPT-3.5.
@@
Code Generation: The FinQA dataset provides solutions in a “program” syntax that, when executed, yields the answer (e.g., in Figure 1 the solution is divide(102400, 619314). However, this derivation does not provide meaningful context of what is being calculated. In our running example, 102400 is not semantically grounded to the document. Koncel-Kedziorski et al. (2023) augments FinQA with readable Python code (including named variables like, dec_shares = 102_400) that can be executed to derive the answer, providing a layer of interpretability. Thus, we use the code-enhanced version of DocFinQA (See Appendix C).
@@
Statistics: The resultant DocFinQA dataset comprises of 5,735 training, 780 development, and 922 test samples, derived from 801 unique SEC filings. Table 1 shows the statistics and characteristics of DocFinQA in comparison with other finance numerical reasoning and long-document QA datasets.
@@
Impact of Data Selection - DocFinQA vs FinQA: Due to the limited availability of complete SEC filings (refer Appendix A) and imperfections in the code generation process, DocFinQA encompasses 7,437 out of 8,281 of FinQA questions. This process may filter out a collection of question types that the LLM did not answer due to its limited capability. We investigates the impact of this process by comparing the distribution of the question types in FinQA and DocFinQA. To do this, we show the distribution of questions grouped by their first 2 non-stop words in Figure 6 (Appendix E). The most important observation is that, overall, the distribution of the question set in DocFinQA and FinQA are very similar. No major groups are being filtered out by our data selection process. The dominant questions (above 1% in FinQA) remain dominant and no major impact on the percentages of those questions is observed. The mid-group (above 0.2% in FinQA) question sets see a mixed effect. A large portion of these questions are increased in percentage while some experience significant loss (e.g., “what percentual” and “what decrease”). Lastly, the long tail group ( under 0.2% in FinQA) either remains the same (e.g., “percent total” and “what greatest”) or is completely wiped out due to a small population (e.g., “was average”, and “what return”).